Find the slope and y-intercept of the given equation.
6x-2y=24
Slope:_______
Y intercept:______
Thanks a bunch!

Answer by nerdybill(7384)

6x-2y=24
Manipulate to "slope-intercept" form:
y = mx + b
-2y = -6x+24
y = 3x-12
.
Slope: 3
Y intercept: -12